Solve the system by substitution.
3x – y = –7
x + y = –9

Answer by checkley71(8403)

x+y=-9 or y=-x-9 now substitute.
3x-(-x-9)=-7
3x+x+9=-7
4x=-7-9
4x=-16
x=-16/4
x=-4 answer.
-4+y=-9
y=-9+4
y=-5 answer.
proof
3*-4-(-5)=-7
-12+5=-7
-7=-7